OpenStudy (anonymous):

Find the exact solutions of the equation that are in the interval [0, 2π ). sin 2t + sin t = 0

OpenStudy (anonymous):

\[\large \sin(2t) + \sin(t) = 0 \implies \sin(2t) = -\sin(t) \implies \sin(2t) = \sin(-t)\] So: general solution is given by: \[\huge 2t = n \pi + (-1)^n \cdot (-t)\] Put n = 0 here: \[\large 2t = -t \implies t = 0\]

OpenStudy (anonymous):

Like wise put n= 1, \[\large 2t = \pi + (-)(-t) \implies 2t = \pi + t \implies t = \pi\]

OpenStudy (anonymous):

Similarly put n= 2: \[\large 2t = 2 \pi - t \implies t = \frac{2 \pi}{3}\]